# https://www.outerlands.io/

### 04 Dec 2023 v1

@@ Abstract
Optimistic rollups are a popular and promising method of increasing the throughput capacity of their underlying chain. These methods rely on economic incentives to guarantee their security. We present a model of optimistic rollups that suggests that the incentives are not necessarily aligned with the expected behavior of the players, thus potentially undermining the security of existing optimistic rollups. ollup, Blockchain, Scaling
@@ 1 Introduction
Optimistic rollups function by having two types of nodes. There are nodes that collect and perform the calculations for the bundle of transactions. They must put up a stake to partacipate. We will call these nodes aggregators, but note that this role is sometimes also called the sequencer. The work done by these nodes is then checked by other nodes that watch the chain and submit fraud proofs when they suspect an inaccuracy. This second role is called a validator. These validators must make a stake before they can press a challenge and they get rewarded if the challenge is successful and penalized if it is not. If a challenge is successful, the validator receives some of the aggregator’s stake and the rest is burned. If a fraud proof is brought, it must be settled on the mainchain, which can be costly, so the system is intended to disincentivize unnecessary proofs [16, 5].
@@ 3 Search costs
A rational, risk neutral $A$ will be indifferent to being honest if $g=1-\frac{s_{A}}{bs_{A}+bf}$.
@@
This is the sum, weighted by $b$, for $V$ going in blind, selecting challenge with probability $g$ and by $(1-b)$ for $V$ performing a search and catching $A$ with certainty.
@@
$h\cdot(-x)+(1-h)\left(\frac{1}{2}s_{A}-x\right)=(1-h)\left(\frac{1}{2}s_{A}\right)-x$
@@
$h=\frac{\frac{1}{2}s_{A}z-zx-s_{A}+(zx+s_{A}x)b}{\frac{1}{2}zs_{A}-s_{A}s_{V}+(z+s_{A})s_{V}b},$ (3)
@@
$\frac{1}{2}s_{A}z-zx-s_{A}+(zx+s_{A}x)b>0$
$\left(\frac{1}{2}s_{A}-x+xb\right)z-s_{A}+s_{A}xb>0$
@@
$(b-1)zx-s_{A}+s_{A}xb<zs_{V}b+(b-1)s_{A}s_{V}$
Dropping the terms without a $z$ yields
$(b-1)zx<zs_{V}b$
given that $10<b<$ 1, the left hand side is always negative and the right hand side can clearly be seen to be positive so the statement holds. Thus there are values for $g$, $h$, and $b$ all strictly between 00 and $1$ that admit an equilibrium. This is not the pure strategy equilibrium from Lemma 2.1 that is intended by the optimistic design. ∎
4 The Role of the Transactor

### 05 Feb 2024 v2

@@ Abstract
Optimistic rollups are a popular and promising method of increasing the throughput capacity of their underlying chain. These methods rely on economic incentives to guarantee their security. We present a model of optimistic rollups that suggests that the incentives are not necessarily aligned with the expected behavior of the players, thus potentially undermining the security of existing optimistic rollups. We discuss some potential solutions illuminated by our model. ollup, Blockchain, Scaling
@@ 1 Introduction
Optimistic rollups function by having two types of nodes. There are nodes that collect and perform the calculations for the bundle of transactions. They must put up a stake to participate. We will call these nodes aggregators, but note that this role is sometimes also called the sequencer. The work done by these nodes is then checked by other nodes that watch the chain and submit fraud proofs when they suspect an inaccuracy. This second role is called a validator. These validators must make a stake before they can press a challenge and they get rewarded if the challenge is successful and penalized if it is not. If a challenge is successful, the validator receives some of the aggregator’s stake and the rest is burned. If a fraud proof is brought, it must be settled on the mainchain, which can be costly, so the system is intended to disincentivize unnecessary proofs [16, 5].
@@ 3 Search costs
A rational, risk neutral $A$ will be indifferent to being honest if $g=1-\frac{s_{A}}{bs_{A}+bz}$.
@@
This is the sum, weighted by $b$, for $V$ going in blind, selecting challenge with probability $g$, and by $(1-b)$ for $V$ performing a search and catching $A$ with certainty.
@@
$h\cdot(-x)+(1-h)\left(\frac{1}{2}s_{A}-x\right)$
which can be simplified to
$(1-h)\left(\frac{1}{2}s_{A}\right)-x$
@@
$h=\frac{\frac{1}{2}s_{A}z-(1-b)(zx+s_{A}x)}{\frac{1}{2}s_{A}z-s_{A}s_{V}+(z+s_{A})s_{V}b},$ (3)
@@
$\frac{1}{2}s_{A}z-zx-s_{A}x+(zx+s_{A}x)b>0$
$\left(\frac{1}{2}s_{A}-x+xb\right)z-s_{A}x+s_{A}xb>0$
@@
$(b-1)(zx+s_{A}x)<zs_{V}b+(b-1)s_{A}s_{V}$
$bzx-zx+bs_{A}x+s_{A}x<zs_{V}b+bs_{A}s_{V}-s_{A}s_{V}$
Once again solving for $b$ yields
$s_{A}s_{V}-s_{A}x-xz<$ $b(zs_{v}+s_{A}s_{V}-xz-s_{A}x)$
$\frac{s_{A}s_{V}-s_{A}x-zx}{s_{A}s_{V}-s_{A}x-zx+zs_{V}}<$ $b$ (4)
We see that this constraint on be conforms to the requirement that $10<b<$ 1. Thus there are values for $g$, $h$, and $b$ all strictly between 00 and $1$ that admit an equilibrium. This is not the pure strategy equilibrium from Lemma 2.1 that is intended by the optimistic design. ∎
3.1 The Role of the Transactor
@@
3.2 Remark
With Theorem 3.1, we have shown a result that has been described before, but we believe our model to be a more thorough articulation of the functioning of optimistic rollups. Our contribution further bolsters the understanding that the verifier’s dilemma poses a significant problem to the well-functioning of optimistic rollups. Of greater consequence, our work also illuminates the roles played by each factor in this problem. For example, if it were not for the validator stake $s_{V}$, or if the validator stake were very small, the lower bound on the probability of no search, $b$, would be $1$ or very close to it, as can be seen in Eq. 4. This would mean that the validator would rather incur the penalty and check randomly than incur the search cost and check with information. Random checking cannot necessarily be written off as a potential solution, as we will discuss in the next section. The precise value of $x$ is also important to the existence and severity of the deviation from the expected functionality. This again points towards a potential solution of alleviating this cost for validators, which we will also discuss in the next section. Furthermore, we can see that there are values for which the equilibrium we have found does not necessarily work. Further study of these values could yield a solution within the existing framework or better illuminate the potential risk of decentralization. All of these observations underscore the importance of the model we present here, which is, to the best of our knowledge, the most accurate and nuanced capture of the optimistic rollup mechanism.
4 Potential Solutions
The verifier’s dilemma has yet to become a practical problem because existing optimistic rollups have yet to decentralize their sequencers. Keeping this status quo is one potential solution, but asking how decentralization might be achieved without compromising the security of the chain is an important question. Another trivial solution would be to revert to a mechanism that requires the validators to check the integrity of the block every time, as is done with proof of stake L1s and zero knowledge (ZK) rollups. As ZK rollups get faster, this may become a more attractive option. We present here the outline of potential solutions that are more in keeping with the spirit of optimistic rollups.
4.1 Random Checks
If we were to stipulate that each block be checked, we would stray away from the spirit of optimistic rollups and lose their lightweight efficiency. Requiring intermittent, random checks could keep the optimistic efficiency without the hazard of the verifier’s dilemma. Whether such a set up would indeed be meaningfully more efficient than simply checking every time would depend on the relative value of the potential benefit from attack and the potential penalties. It would also depend on precisely how the random checks would be implemented. This would not necessitate fraud proofs, which are needed to find precisely where the chain became faulty, and instead would require proofs of just the block in question. These proofs could be done by nodes, which would again need incentives, or by smart contracts on the L1, which could have expensive execution that might fall on the users or be paid for with inflation. If we assume this could be done with a smart contract without too much additional cost to the user, we get a game where $\mathcal{C}$ is the contract ‘player’ in the place of $V$. Let us also define $p$, where $10<p<$ 1, to be the probability with which the contract checks the proof.
Lemma 5
There are values of the probability of check, $p$, $s_{A}$ and $z$ such that a rational, risk neutral aggregator will always be honest.
{istgame}
[…]
Game 3: Game tree illustrating a random check where utility is that of $A$ and the dotted line indicates an information set.
It is immediate that the expected income on the left child of the root is zero. Recall this is because we normalized it so. Since the left, honest choice on the part of $A$ is the desired outcome, we need a $p$ such that the right side yields negative income in expectation. Thus we would need,
$(1-p)z-ps_{A}<0$
$(1-p)z<ps_{A}$
Solving for $p$ yields,
$ps_{A}>z-pz$
$p(s_{A}+z)>z$
$p>\frac{z}{z+s_{A}}$
This is clearly in keeping with our requirement that $10<p<$ 1, but is also clear that, given $z>s_{A}$ and especially in cases where $z$ is much larger than $s_{A}$, $p$ will be close to $1$. Thus, there are only minor savings as compared to checking every time, but there are values for which such a setup would result in a rational, risk neutral $A$ behaving honestly, as desired. ∎
In this proof, we have assumed that there are only random checkers and no traditional validators. This is a viable solution, although it requires the approximate value of $z$ to be known and for the smart contract to work as described. A scenario with both random and traditional nodes might yield a different, potentially more efficient, result.
4.2 Random Mistakes
The motivating story for the verifier’s dilemma is a scenario in which the aggregator makes no mistakes and thus removes the incentive to ever check. This in turn leads to an opening for mistakes or malice on the part of the aggregator. Adding mistakes by design thus seems like a potential way to incentivize watchfulness on the part of the validators. This might pose a practical problem as presumably only some of the aggregator nodes, say those belonging to the foundation associated with the chain, would be making these purposeful mistakes and validators might then only check the blocks proposed by those nodes. Furthermore some entity would have to incur the cost of the stakes slashed when the mistakes are caught.
As we have shown in the proof for Eq. 2, there is an expression for $h$ for which the validator is indifferent between search and no search. Thus, we can presume that there are still values of $b$ and $g$ between 00 and $1$ that will yield a mixed equilibrium different from that desired. Modeling this precisely would require coefficients for both whether the block is being produced by a mistake-making $A$ and, if so, whether it elects to make a mistake. A non-mistake-making $A$ would have to take these terms into account. We leave the realization of such a model to future work, but it is far from immediate that reasonable values would yield an equilibrium in which the validator searches every time. It would likely be unreasonable for a mistake-making $A$ to make mistakes every time, as this would be tremendously expensive.
4.3 Easter Eggs
If the search costs, $x$, are relatively small, compensating the validators directly for their services seems like a tempting solution. However, it may not be clear whether or not the search is being performed correctly and checking this search would be yet another role with its own incentives. If it were possible to hide reward ‘Easter eggs,’ which could be found through proper searching but not a random challenge, this could act as a delivery mechanism for validator rewards. These rewards would be realized even when the aggregator behaves honestly.
Lemma 6
(Sketch) Compensating the validator with some term $e\geq x$, will result in the expected income for the validator of always searching being higher than not searching.
Let us add $e$, the expected value of finding the Easter egg, to the validator income in the case of search. Reading off from the proof of Eq. 2, we have that search yields more income if
$(1-h)\frac{1}{2}s_{A}-x+e>-ghs_{V}+g(1-h)\frac{1}{2}$
since $g<1$, it is clear to see that if we have $-x+e>-ghs_{V}$, the above term is always true. If we require that $e\geq x$, the result is immediate. ∎